# Direct comparison of stochastic driven nonlinear dynamical systems for combinatorial optimization

### v2

@@ I Introduction
More recently, various new computing paradigms have been proposed and realized in different physical systems to more effectively traverse energy landscapes by allowing the original binary variable $\sigma_{l}=\pm 1$ to relax to a continuous variable of either real $x_{l}\in\mathbb{R}$ or complex $\Psi_{l}\in\mathbb{C}$ type. In these algorithms, optimization is achieved by mapping local minima to fixed points through bifurcation processes which are usually created by tuning a set of gain/dissipative strengths (or equivalently, proper driven effects) in the system Strogatz (2015). Throughout this work, we refer to these physical systems as stochastic driven nonlinear dynamical systems (SDNDSs), because the name highlights their three common key features, namely, stochastic noise terms, driven or gain/dissipative effects, and nonlinearity.
@@
Furthermore, we conduct a range of experiments to deeper investigate the performance of these heuristics. First, because the dynamical systems studied here are simulated by numerically solving the associated equations of motion (EoM) which are usually ordinary differential equations (ODEs), we study how using different numerical integration methods—e.g., Runge–Kutta (RK)—affect performance. To better understand how graph connectivity affects performance, we systematically dilute a fully-connected Sherrington-Kirkpatrick model Sherrington and Kirkpatrick (1975) to a Viana-Bray model Viana and Bray (1985). To this end, we compare the performance of SA against the SBM, similar to the recent work presented in Ref. Pawlowski et al. (2025).
Table 1: List of different stochastic driven nonlinear dynamical systems studied in this work. The artificial spin is denoted by either $x_{l}$ (real field) or $\Psi_{l}$ (complex field). Other common parameters include the pumping/driving strength $\alpha$, global Ising coupling $\beta$, and the nonlinear activation function $\phi(\bullet)$. For other model-specific parameters, please refer to the relevant subsections in Sec. II.
@@
GD $\begin{cases}\dot{\Psi}_{l}=\Psi_{l}\left(\alpha_{l}-\alpha_{c0}-|\Psi_{l}|^{2}\right)\\ \hskip 22.76228pt+\sum\limits_{m}\beta_{lm}K_{lm}\Psi_{m}+\sum\limits_{q=1}^{2}h_{ql}\Psi_{l}^{*}+g_{l},\\ \dot{\alpha}_{l}=\epsilon\left(\rho_{\text{th}}-|\Psi_{l}|^{2}\right),\vspace{1mm}\\ \dot{K}_{lm}=\epsilon^{\prime}\left(J_{lm}-\beta_{lm}K_{lm}\right)\end{cases}$
SBM $\begin{cases}\dot{x}_{l}=\Delta y_{l},\vspace{2mm}\\ \dot{y}_{l}=-\left(Kx_{l}^{2}-\alpha(t)+\Delta\right)x_{l}+\beta\sum\limits_{m}J_{lm}x_{m}\end{cases}$
@@ II Overview of physical models
In this section, we give a brief overview of the SDNDSs studied. We present the underlying physics of these algorithms and make comparisons between them. Besides the recently-proposed SDNDSs, such as different Ising machines and gain-dissipative methods, we also discuss the Hopfield neural network, which is intrinsically a nonlinear dynamical system. For the convenience of the reader, we have summarized all the SDNDSs studied here in Table. 1. Throughout this work, we adopt a consistent notation for the corresponding EoMs.
@@ II.1 Coherent Ising machine and its variations
Physical Ising machines usually suffers from a limited number of artificial spins when operating in the quantum regime. On the other hand, a classical system like an optical platform has been used to demonstrate the so-called coherent Ising machine (CIM), which can tackle optimization problems with thousands of nodes Inagaki et al. (2016b). The CIM architecture is primarily based on an optical parametric oscillator (OPO) consisting of a fiber loop and a degenerate parametric amplifier or a single-mode squeezer. The OPO is pumped above its oscillation threshold, hence exciting the pulse modes circulating inside. Thanks to the phase sensitivity of the optical squeezer, the phases of these modes tend to become locked at either 00 or $\pi$, giving rise to a binary constraint.
@@ II.1.1 Simulated coherent Ising machine
where $x_{\text{sat}}$ defines a hard-wall boundary and confines the spins within the interval $[-x_{\text{sat}},x_{\text{sat}}]$.
@@ II.1.2 Optoelectronic oscillator coherent Ising machine
In this dynamical system, a rather unique activation function, i.e., $\cos^{2}(\bullet)$ appears. This is justified by the Taylor expansion near $x_{l}=0$ because the spin amplitude is small when the system is driven to the laser threshold ($\alpha\approx 1$) where CIMs have their optimal performance Leleu et al. (2017). Assuming that the pumping term is dominant over the Ising potential, namely, $\alpha x_{l}\gg\beta\sum_{m}J_{lm}x_{m}$ and truncating the corresponding Taylor series at the third order, we can easily recover the equations of motion of the OPO-CIM from those of the OEO-CIM (see Table. 1). The Sim-CIM is also similar to OEO-CIM except with a different activation function.
@@ II.2 Gain-dissipative platforms
The coefficients $\beta_{lm}K_{lm}$ represent the coupling strength between $\Psi_{l}$ and $\Psi_{m}$ in which $\beta_{lm}$ arises from particle injection effects, whereas $K_{lm}$ is induced by other mechanisms like bare hopping or the laser-controlled coupling Berloff et al. (2017). Unlike the global Ising coupling $\beta$ in other models, GD systems use a tunable local coupling strength $\beta_{lm}$ instead (see Table. 1). Physically, a common choice for the coupling coefficients is $\beta_{lm}=\alpha_{l}+\alpha_{m}$, where the onsite injection rate $\alpha_{l}$ is governed by the following dynamics:
@@
Unlike the two CIM models discussed earlier, the GD platform is highly versatile, in the sense that it can be tuned to suit different types of optimization problems. For example, if one sets $h_{ql}=0$ in Eq. (7), the GD model can be used to optimize the XY Hamiltonian Yeomans (1992). On the other hand if $h_{2m}>\sum_{l\neq m}|J_{lm}|$, the phases will be fixed at 00 or $\pi$, and therefore, a binary variable problem can be simulated. With other values of $h_{ql}$, the algorithm can fit a $q$-state clock model in which the phases are forced to be $2\pi n/q,n\in\mathbb{Z}$ Kalinin and Berloff (2018b). The GD system’s adaptability can also be exploited in other nontrivial ways. For instance, a recent work demonstrated a physics-inspired algorithm that solves higher-order polynomial binary optimization problems without mapping them to a quadratic unconstrained binary form Stroev and Berloff (2021). However, higher-order terms become more difficult to numerically integrate and might lead to numerical instabilities. As another example, by breaking the $U(1)$ symmetry of the order parameter into $\mathbb{Z}_{3}$, a ternary spin system can be introduced to tackle the Max-3-SAT problem Harrison et al. (2020) natively.
@@ II.4 Simulated bifurcation machine
Here, $K$ denotes the inter-oscillator interaction strength, which eventually becomes an onsite but site-dependent potential. A similar phenomenon is realized in the GD model via the inter-atomic interaction $-|\Psi_{l}|^{2}$ and through the nonlinear loss function $-x_{l}^{3}$ in the OPO-CIM. In the Sim-CIM or OEO-CIM, the effect of such a term is implicitly replaced by a nonlinear activation function. Physically, a strong density interaction prevents any extra population from accumulating in a single state. Similarly, the activation function bounds the variables to be in $[-1,1]$.
@@ III.1 Methodology
The Ising test sets are tile-planted problems Hamze et al. (2018); Perera et al. (2020) generated by using the Chook Perera et al. (2021) benchmark instance generator. These problems consist of square lattices of linear size $L$ with periodic boundary conditions in both directions and $N=L^{2}$ spins. Each lattice site is connected to four neighboring sites with a coupling strength that can only be ±$\pm 1$ or ±$\pm 2$. We also study the Viana-Bray Viana and Bray (1985) model with variable connectivity in Sec. IV.2 to better understand if graph connectivity influences the algorithmic performance.
@@
Figure 1: Time-to-solution (TTS) of various SDNDSs listed in Table. 1 versus the linear size of the two-dimensional lattice. The dashed curves represent linear fits in the logarithmic (base 10) scale to the five largest system sizes.
@@ III.2 Benchmarking results
Table 2: Optimal parameters for the Sim-CIM simulations for different linear problem sizes $L$. $\alpha_{0}$ and $\alpha_{1}$ determine the pumping profile (hyperbolic in this case), $\beta$ is the global Ising coupling strength, and $x_{\text{sat}}$ is the spin saturation value. In addition, $\sigma_{g}$ is the width of the Gaussian noise function and $N_{t}$ is the total number of iterations which is held fixed here.
@@
$4$ $0.7$ $1$ $0.2$ $3$ $0.1$ $1000$
$6$ $0.7$ $1$ $0.2$ $3$ $0.1$ $1000$
$8$ $0.6$ $1.5$ $0.2$ $3$ $0.1$ $1000$
$10$ $0.6$ $1.5$ $0.2$ $3$ $0.1$ $1000$
$12$ $0.6$ $1.5$ $0.2$ $3$ $0.1$ $1000$
$14$ $0.6$ $1.5$ $0.2$ $3$ $0.1$ $1000$
$16$ $0.6$ $1.5$ $0.2$ $5$ $0.2$ $1000$
$18$ $0.6$ $2$ $0.2$ $5$ $0.2$ $1000$
$20$ $0.6$ $2.5$ $0.2$ $5$ $0.2$ $1000$
Simulate coherent Ising machine.— For the Sim-CIM, there are several tunable parameters including the hyperbolic pumping profile (sharpness $\alpha_{1}$ and maximum pumping strength $\alpha_{0}$), the global Ising coupling strength $\beta$, and the spin saturation value $x_{\text{sat}}$. Given the problem sizes studied in this work, we use a constant total number of iterations of $N_{t}=1000$. We present the optimized parameters of the Sim-CIM algorithm in Tab. 2 and the corresponding TTS scaling in Fig. 1. Due to the fixed number of steps, the Sim-CIM has a larger TTS for smaller problems, but it becomes competitive as the problem size increases. Our experiments suggest that the performance of the Sim-CIM is sensitive to parameter choices. Although $\alpha_{0}$ and $\beta$ are relatively robust, $\alpha_{1}$, $x_{\text{sat}}$, and $\sigma_{g}$ vary depending on the size of the problem.
Optoelectronic oscillator coherent Ising machine.— The OEO-CIM has three tunable parameters, namely, the gain strength $\alpha$, the global Ising coupling $\beta$, and the noise level $\sigma_{g}$. In Ref. Böhm et al. (2019), constants $\alpha=0.25$ and $\beta=0.29$ are chosen for square lattice problems. These values are consistent with our optimized results reported in Tab. 3. Because the OEO-CIM converges relatively fast in comparison to the Sim-CIM—usually in a few hundred steps—it is very competitive for small system sizes. However, the algorithm becomes unstable with larger system sizes, leading to unreasonably large errors in the TTS when $L>10$. Thus, the benchmark results for the OEO-CIM in Fig. 1 are obtained by introducing a time-dependent gain profile. Table 3 shows that the optimal parameters of the OEO-CIM follow a trend similar to that of the parameters of the Sim-CIM. However, we observe a sudden jump in the gain sharpness $\alpha_{1}$ at $L=18$, beyond which the OEO-CIM can rarely find the ground state after $100$ restarts for any of the $100$ instances studied.
Table 3: Optimal Parameters for the OEO-CIM simulations for different linear problem sizes $L$. $\alpha_{0}$ and $\alpha_{1}$ define a hyperbolic pumping profile, $\beta$ is the global Ising coupling strength, $\sigma_{g}$ is the Gaussian noise level, and finally $N_{t}$ is the total number of iterations.
@@
$4$ $0.45$ $2$ $0.29$ $0.1$ $2^{6}$
$6$ $0.45$ $3.5$ $0.29$ $0.1$ $2^{6}$
$8$ $0.45$ $5$ $0.29$ $0.1$ $2^{7}$
$10$ $0.45$ $5$ $0.29$ $0.1$ $2^{7}$
$12$ $0.45$ $5$ $0.29$ $0.2$ $2^{8}$
$14$ $0.45$ $5$ $0.29$ $0.2$ $2^{8}$
$16$ $0.45$ $5$ $0.29$ $0.2$ $2^{9}$
$18$ $0.45$ $9$ $0.29$ $0.2$ $2^{9}$
Gain dissipative algorithm We choose the parameters of the GD algorithm following the original proposal of Ref. Kalinin and Berloff (2018a). Specifically, we use a uniform local Ising coupling $\beta_{lm}=\beta=1$ and $K_{lm}=J_{lm}$ such that only two of three equations of motion are nontrivial (see Tab. 1). The remaining parameters are chosen as
@@
and $\Psi_{i}(t=0)=1$. Finally, $h_{2m}$ has a hyperbolic tangent profile specified by $\alpha_{1}$ and $\alpha_{0}$. In our experiments, we find that, unlike CIMs, GD systems can be robust in terms of parameter choices, which is consistent with the authors’ results on Max-Cut problems reported in Ref. Kalinin and Berloff (2018a). Thus, we have fixed the gain profile and the global Ising coupling for different problem sizes and only tuned the total number of iterations. The parameters used in the simulations are summarized in Tab. 4. Figure 1 shows that our implementation of the GD algorithm has a relatively large TTS for all problem sizes studied. We conjecture that the synchronization of the network may take extra time due to the small pumping rate in our experiments and because numerical computations using complex numbers are slower than when using real numbers.
Table 4: Total number of iterations $N_{t}$ for different dynamical system algorithms versus linear problem sizes $L$ used in the simulations. For simulated annealing (SA), we use an inverse linear annealing schedule and set the final inverse temperature to $\beta_{f}=3$. For the GD system we fix $\alpha_{0}=0.5$, $\alpha_{1}=10$, $\beta=1$, and $\sigma_{g}=0.2$. For the SBM, the time step is fixed to $dt=0.5$. In the case of the Hopfield model dynamical system, we set $\zeta_{0}=10$, $\zeta_{d}=0.8$, $dt=0.5$, and $\sigma_{g}=0.1$. For the HNN model with fourth-order Runge-Kutta (RK4) and Runge-Kutta-Fehlberg (RK45) integration schemes, we fix the initial time step to $dt=1$ and $dt_{0}=1$, respectively. We also choose $\epsilon=1$ in order to curb errors.
$L$ $4$ $6$ $8$ $10$ $12$ $14$ $16$ $18$ $20$

### v3

# Direct comparison of stochastic-driven nonlinear dynamical systems for combinatorial optimization

@@ I Introduction
More recently, various new computing paradigms have been proposed and realized in different physical systems to more effectively traverse energy landscapes by allowing the original binary variable $\sigma_{l}=\pm 1$ to relax to a continuous variable of either real $x_{l}\in\mathbb{R}$ or complex $\Psi_{l}\in\mathbb{C}$ type. In these algorithms, optimization is achieved by mapping local minima to fixed points through bifurcation processes which are usually created by tuning a set of gain/dissipative strengths (or equivalently, proper driven effects) in the system Strogatz (2015). Throughout this work, we refer to these physical systems as stochastic-driven nonlinear dynamical systems (SDNDSs), because the name highlights their three common key features, namely, stochastic noise terms, driven or gain/dissipative effects, and nonlinearity.
@@
Furthermore, we conduct a range of experiments to deeper investigate the performance of these heuristics. First, because the dynamical systems studied here are simulated by numerically solving the associated equations of motion (EoM) which are usually ordinary differential equations (ODEs), we study how using different numerical integration methods—e.g., Runge–Kutta (RK)—affect performance. To better understand how graph connectivity affects performance, we systematically dilute a fully-connected Sherrington-Kirkpatrick model Sherrington and Kirkpatrick (1975) to a Viana-Bray model Viana and Bray (1985). To this end, we compare the performance of SA against the SBM.
Table 1: List of different stochastic-driven nonlinear dynamical systems studied in this work. The artificial spin is denoted by either $x_{l}$ (real field) or $\Psi_{l}$ (complex field). Other common parameters include the pumping/driving strength $\alpha$, global Ising coupling $\beta$, and the nonlinear activation function $\phi(\bullet)$. For other model-specific parameters, please refer to the relevant subsections in Sec. II.
@@
GD $\begin{cases}\dot{\Psi}_{l}=\Psi_{l}\left(\alpha_{l}-\alpha_{c0}-|\Psi_{l}|^{2}\right)\\ \hskip 22.76228pt+\sum\limits_{m}\beta_{lm}K_{lm}\Psi_{m}+\sum\limits_{q=1}^{2}h_{ql}\Psi_{l}^{*}+g_{l},\\ \dot{\alpha}_{l}=\epsilon\left(\rho_{\text{th}}-|\Psi_{l}|^{2}\right),\vskip 2.84526pt\\ \dot{K}_{lm}=\epsilon^{\prime}\left(J_{lm}-\beta_{lm}K_{lm}\right)\end{cases}$
SBM $\begin{cases}\dot{x}_{l}=\Delta y_{l},\vskip 5.69054pt\\ \dot{y}_{l}=-\left(Kx_{l}^{2}-\alpha(t)+\Delta\right)x_{l}+\beta\sum\limits_{m}J_{lm}x_{m}\end{cases}$
@@ II Overview of physical models
In this section, we give a brief overview of the SDNDSs studied. We present the underlying physics of these algorithms and make comparisons between them. Besides the recently proposed SDNDSs, such as different Ising machines and gain-dissipative methods, we also discuss the Hopfield neural network, which is intrinsically a nonlinear dynamical system. For the convenience of the reader, we have summarized all the SDNDSs studied here in Table 1. Throughout this work, we adopt a consistent notation for the corresponding EoMs.
@@ II.1 Coherent Ising machine and its variations
Physical Ising machines usually suffers from a limited number of artificial spins when operating in the quantum regime. However, a classical system like an optical platform has been used to demonstrate the so-called coherent Ising machine (CIM), which can tackle optimization problems with thousands of nodes Inagaki et al. (2016b). The CIM architecture is primarily based on an optical parametric oscillator (OPO) consisting of a fiber loop and a degenerate parametric amplifier or a single-mode squeezer. The OPO is pumped above its oscillation threshold, hence exciting the pulse modes circulating inside. Thanks to the phase sensitivity of the optical squeezer, the phases of these modes tend to become locked at either 0 or $\pi$, giving rise to a binary constraint.
@@ II.1.1 Simulated coherent Ising machine
where $x_{\text{sat}}$ defines a hard-wall boundary and confines the spins within the interval $x_{\text{sat}},x_{\text{sat}}]$.
@@ II.1.2 Optoelectronic oscillator coherent Ising machine
In this dynamical system, a rather unique activation function, i.e., $\cos^{2}(\bullet)$ appears. This is justified by the Taylor expansion near $x_{l}=0$ because the spin amplitude is small when the system is driven to the laser threshold ($\alpha\approx 1$) where CIMs have their optimal performance Leleu et al. (2017). Assuming that the pumping term is dominant over the Ising potential, namely, $\alpha x_{l}\gg\beta\sum_{m}J_{lm}x_{m}$ and truncating the corresponding Taylor series at the third order, we can easily recover the equations of motion of the OPO-CIM from those of the OEO-CIM (see Table 1). The Sim-CIM is also similar to OEO-CIM except with a different activation function.
@@ II.2 Gain-dissipative platforms
The coefficients $\beta_{lm}K_{lm}$ represent the coupling strength between $\Psi_{l}$ and $\Psi_{m}$ in which $\beta_{lm}$ arises from particle injection effects, whereas $K_{lm}$ is induced by other mechanisms like bare hopping or the laser-controlled coupling Berloff et al. (2017). Unlike the global Ising coupling $\beta$ in other models, GD systems use a tunable local coupling strength $\beta_{lm}$ instead (see Table 1). Physically, a common choice for the coupling coefficients is $\beta_{lm}=\alpha_{l}+\alpha_{m}$, where the onsite injection rate $\alpha_{l}$ is governed by the following dynamics:
@@
Unlike the two CIM models discussed earlier, the GD platform is highly versatile, in the sense that it can be tuned to suit different types of optimization problems. For example, if one sets $h_{ql}=0$ in Eq. (7), the GD model can be used to optimize the XY Hamiltonian Yeomans (1992). However, if $h_{2m}>\sum_{l\neq m}|J_{lm}|$, the phases will be fixed at 0 or $\pi$, and therefore, a binary variable problem can be simulated. With other values of $h_{ql}$, the algorithm can fit a $q$-state clock model in which the phases are forced to be $2\pi n/q,n\in\mathbb{Z}$ Kalinin and Berloff (2018b). The GD system’s adaptability can also be exploited in other nontrivial ways. For instance, a recent work demonstrated a physics-inspired algorithm that solves higher-order polynomial binary optimization problems without mapping them to a quadratic unconstrained binary form Stroev and Berloff (2021). However, higher-order terms become more difficult to numerically integrate and might lead to numerical instabilities. As another example, by breaking the $U(1)$ symmetry of the order parameter into $\mathbb{Z}_{3}$, a ternary spin system can be introduced to tackle the Max-3-SAT problem Harrison et al. (2020) natively.
@@ II.4 Simulated bifurcation machine
Here, $K$ denotes the inter-oscillator interaction strength, which eventually becomes an onsite but site-dependent potential. A similar phenomenon is realized in the GD model via the inter-atomic interaction $|\Psi_{l}|^{2}$ and through the nonlinear loss function $x_{l}^{3}$ in the OPO-CIM. In the Sim-CIM or OEO-CIM, the effect of such a term is implicitly replaced by a nonlinear activation function. Physically, a strong density interaction prevents any extra population from accumulating in a single state. Similarly, the activation function bounds the variables to be in [1,1][-1,1].
@@ III.1 Methodology
The Ising test sets are tile-planted problems Hamze et al. (2018); Perera et al. (2020) generated by using the Chook Perera et al. (2021) benchmark instance generator. These problems consist of square lattices of linear size $L$ with periodic boundary conditions in both directions and $N=L^{2}$ spins. Each lattice site is connected to four neighboring sites with a coupling strength that can only be $\pm 1$ or $\pm 2$. We also study the Viana-Bray Viana and Bray (1985) model with variable connectivity in Sec. IV.2 to better understand if graph connectivity influences the algorithmic performance.
@@
Figure 1: Time-to-solution (TTS) of various SDNDSs listed in Table 1 versus the linear size of the two-dimensional lattice. The dashed curves represent linear fits in the logarithmic (base 10) scale to the five largest system sizes.
@@ III.2 Benchmarking results
Table 2: Optimal parameters for the Sim-CIM simulations for different linear problem sizes $L$. $\alpha_{0}$ and $\alpha_{1}$ determine the pumping profile (hyperbolic in this case), $\beta$ is the global Ising coupling strength, $x_{\text{sat}}$ is the spin saturation value, $\sigma_{g}$ is the width of the Gaussian noise function, and $N_{t}$ is the total number of iterations which is held fixed here.
@@
44 $0.7$ 11 $0.2$ 33 $0.1$ $1000$
66 $0.7$ 11 $0.2$ 33 $0.1$ $1000$
88 $0.6$ $1.5$ $0.2$ 33 $0.1$ $1000$
$10$ $0.6$ $1.5$ $0.2$ 33 $0.1$ $1000$
$12$ $0.6$ $1.5$ $0.2$ 33 $0.1$ $1000$
$14$ $0.6$ $1.5$ $0.2$ 33 $0.1$ $1000$
$16$ $0.6$ $1.5$ $0.2$ 55 $0.2$ $1000$
$18$ $0.6$ 22 $0.2$ 55 $0.2$ $1000$
2020 $0.6$ $2.5$ $0.2$ 55 $0.2$ $1000$
III.2.1 Simulated coherent Ising machine
For the Sim-CIM, there are several tunable parameters including the hyperbolic pumping profile (sharpness $\alpha_{1}$ and maximum pumping strength $\alpha_{0}$), the global Ising coupling strength $\beta$, and the spin saturation value $x_{\text{sat}}$. Given the problem sizes studied in this work, we use a constant total number of iterations of $N_{t}=1000$. We present the optimized parameters of the Sim-CIM algorithm in Table 2 and the corresponding TTS scaling in Fig. 1. Due to the fixed number of steps, the Sim-CIM has a larger TTS for smaller problems, but it becomes competitive as the problem size increases. Our experiments suggest that the performance of the Sim-CIM is sensitive to parameter choices. Although $\alpha_{0}$ and $\beta$ are relatively robust, $\alpha_{1}$, $x_{\text{sat}}$, and $\sigma_{g}$ vary depending on the size of the problem.
III.2.2 Optoelectronic oscillator coherent Ising machine
The OEO-CIM has three tunable parameters, namely, the gain strength $\alpha$, the global Ising coupling $\beta$, and the noise level $\sigma_{g}$. In Ref. Böhm et al. (2019), constants $\alpha=0.25$ and $\beta=0.29$ are chosen for square lattice problems. These values are consistent with our optimized results reported in Table 3. Because the OEO-CIM converges relatively fast in comparison to the Sim-CIM—usually in a few hundred steps—it is very competitive for small system sizes. However, the algorithm becomes unstable with larger system sizes, leading to unreasonably large errors in the TTS when $L>10$. Thus, the benchmark results for the OEO-CIM in Fig. 1 are obtained by introducing a time-dependent gain profile. Table 3 shows that the optimal parameters of the OEO-CIM follow a trend similar to that of the parameters of the Sim-CIM. However, we observe a sudden jump in the gain sharpness $\alpha_{1}$ at $L=18$, beyond which the OEO-CIM can rarely find the ground state after $100$ restarts for any of the $100$ instances studied.
Table 3: Optimal parameters for the OEO-CIM simulations for different linear problem sizes $L$. $\alpha_{0}$ and $\alpha_{1}$ define a hyperbolic pumping profile, $\beta$ is the global Ising coupling strength, $\sigma_{g}$ is the Gaussian noise level, and $N_{t}$ is the total number of iterations.
@@ III.2.2 Optoelectronic oscillator coherent Ising machine
44 $0.45$ 22 $0.29$ $0.1$ $2^{6}$
66 $0.45$ $3.5$ $0.29$ $0.1$ $2^{6}$
88 $0.45$ 55 $0.29$ $0.1$ $2^{7}$
$10$ $0.45$ 55 $0.29$ $0.1$ $2^{7}$
$12$ $0.45$ 55 $0.29$ $0.2$ $2^{8}$
$14$ $0.45$ 55 $0.29$ $0.2$ $2^{8}$
$16$ $0.45$ 55 $0.29$ $0.2$ $2^{9}$
$18$ $0.45$ 99 $0.29$ $0.2$ $2^{9}$
III.2.3 Gain dissipative algorithm
We choose the parameters of the GD algorithm following the original proposal of Ref. Kalinin and Berloff (2018a). Specifically, we use a uniform local Ising coupling $\beta_{lm}=\beta=1$ and $K_{lm}=J_{lm}$ such that only two of three equations of motion are nontrivial (see Table 1). The remaining parameters are chosen as
@@ III.2.3 Gain dissipative algorithm
and $\Psi_{i}(t=0)=1$. Finally, $h_{2m}$ has a hyperbolic tangent profile specified by $\alpha_{1}$ and $\alpha_{0}$. In our experiments, we find that, unlike CIMs, GD systems can be robust in terms of parameter choices, which is consistent with the authors’ results on Max-Cut problems reported in Ref. Kalinin and Berloff (2018a). Thus, we have fixed the gain profile and the global Ising coupling for different problem sizes and only tuned the total number of iterations. The parameters used in the simulations are summarized in Table 4. Figure 1 shows that our implementation of the GD algorithm has a relatively large TTS for all problem sizes studied. We conjecture that the synchronization of the network may take extra time due to the small pumping rate in our experiments and because numerical computations using complex numbers are slower than when using real numbers.
Table 4: Total number of iterations $N_{t}$ for different dynamical system algorithms versus linear problem sizes $L$ used in the simulations. For simulated annealing (SA), we use an inverse linear annealing schedule and set the final inverse temperature to $\beta_{f}=3$. For the GD system we fix $\alpha_{0}=0.5$, $\alpha_{1}=10$, $\beta=1$, and $\sigma_{g}=0.2$. For the SBM, the time step is fixed to $dt=0.5$. In the case of the Hopfield model dynamical system, we set $\zeta_{0}=10$, $\zeta_{d}=0.8$, $dt=0.5$, and $\sigma_{g}=0.1$. For the HNN model with fourth-order Runge-Kutta (RK4) and Runge-Kutta-Fehlberg (RK45) integration schemes, we fix the initial time step to $dt=1$ and $dt_{0}=1$, respectively. We also choose $\epsilon=1$ to curb errors.
$L$ 44 66 88 $10$ $12$ $14$ $16$ $18$ 2020